Solving by elimination..
3x+2y=9
-x+3y=8

Answer by checkley71(8403)

3X+2Y=9 NOW MULTIPLY THE SECOND EQUATION BY 3 TO GET
-3X+9Y=24 NOW ADD THESE TWO EQUATIONS
--------------
+11Y=33 OR Y=33/11 OR Y=3 THEN 3X+2*3=9 OR 3X+6=9 OR 3X=3 OR X=1
PROOF -1*1+3*3=8 OR -1+9=8 OR 8=8